### v1

# Hallmarks of Spin Textures for High-Harmonic Generation

Abstract
Spin-orbit coupling and quantum geometry are fundamental aspects in modern condensed matter physics, with their primary manifestations in momentum space being spin textures and Berry curvature. In this work, we investigate their interplay with high-harmonic generation (HHG) in two-dimensional non-centrosymmetric materials, with an emphasis on even-order harmonics. Our analysis reveals that the emergence of finite even-order harmonics necessarily requires a broken twofold rotational symmetry in the spin texture, as well as a non-trivial Berry curvature in systems with time-reversal invariance. This symmetry breaking can arise across various degrees of freedom and impact both spin textures and optical response via spin-orbit interactions. These findings underscore the potential of HHG as a powerful tool for exploring electronic phases with broken rotational symmetry, as well as the associated phase transitions in two-dimensional materials. This approach provides novel perspectives for designing symmetry-dependent nonlinear optical phenomena.
Introduction — Recent technological advances in laser science have enabled the development of high-intensity light sources across a wide range of energies. Such strong pulses have been increasingly employed in experimental condensed matter physics to induce nonlinear optical responses in crystalline solids. A notable example is high-harmonic generation (HHG), a nonlinear process in which a material exposed to light at a specific frequency emits radiation at multiples of this fundamental frequency[1, 2]. A key distinction in solid-state HHG is that odd-order harmonics are universally present, whereas even-order harmonics are suppressed by inversion symmetry. Surface, interfaces and certain two-dimensional (2D) materials where inversion symmetry is naturally broken offer potential avenues for observing even-order harmonic emission. 2D non-centrosymmetric structures of materials with substantial spin-orbit coupling (SOC) are generally characterized by a finite Rashba SOC[3, 4]. This leads to energy band splitting and the emergence of momentum-dependent spin textures – a variation of the spin direction over the Brillouin zone. At interfaces with trigonal crystal symmetries, these spin textures can also induce a non-zero Berry curvature even if magnetic order is absent[5, 6, 7]. The presence of SOC and Berry curvature in two dimensions is of significant importance for spintronics[8, 9, 10] and topological phenomena[11], respectively. Moreover, recent studies have shown that both SOC and Berry curvature are fundamental physical ingredients for HHG: the former governs terahertz third-harmonic generation in transition metals[12], while a direct correspondence between the latter and second-harmonic generation has been established in spin-orbit-free systems[13, 14, 15, 16].
A relevant question is whether and how Berry curvature and SOC cooperate in non-centrosymmetric 2D materials for HHG, particularly in the context of even-order harmonics. Recent studies on the nonlinear Hall effect with time-reversal symmetry[17, 7, 18, 19, 20] have established a close connection between SOC, Berry curvature, and second-harmonic generation using a semiclassical approach[21]. The aim of this Letter is to demonstrate, using HHG selection rules[22] and explicit calculations, the interplay between SOC, quantum geometry, and high-order harmonics in 2D non-centrosymmetric systems, with a focus on the relation between the emergence of non-zero even-order harmonics and distinct symmetry-allowed spin textures. Note that in certain magnetic materials, non-zero second-order response has been attributed to distinctive magnetic patterns both with[23, 24] and without[25, 26] SOC, but without accounting for the contribution of the Berry curvature and the symmetry properties of the spin texture. Our analysis establishes that a spin texture that breaks a $\hat{\mathcal{C}}_{2}$ twofold rotational symmetry is necessary for observing even-order harmonics. The Berry curvature imposes constraints only in time-reversal symmetric conditions: its non-vanishing is essential for the appearance of even-order harmonics. By contrast, when time-reversal symmetry is broken even-order harmonics can persist even in the complete absence of Berry curvature.
[…]
${\bf J}(t)=\sum_{n}{\bf J}_{n}e^{in\Omega t},$ (1)
Figure 1: Illustration of the patterns of the spin-texture in momentum space for the high-harmonics generation. Left panel: in 2D systems with spin texture invariant under twofold rotational symmetry, $\hat{\mathcal{C}}_{2}$, there is no emission of even-order harmonics. Right panel: for a spin texture exhibiting a $\hat{\mathcal{C}}_{2}$ symmetry breaking, the emission of even-order harmonics is allowed.
where the Fourier component ${\bf J}_{n}$ encodes the information on the emission of the $n$-th-order harmonic. The time-dependent Hamiltonian $\hat{H}(t)$ is generally not invariant under the static point-group symmetry $\hat{\mathcal{P}}$. As we show in details in [28], one must find a suitable combination of the point-group symmetry and an operator acting on the time variable. This combined symmetry, referred to as dynamical symmetry[29, 30, 31, 22], must ensure that both the momentum and the vector potential transform identically. Consider, for instance, the $\hat{\mathcal{C}}_{2}$ symmetry operation, which rotates the system by $\pi$ around the $z$ axis, resulting in a sign flip of both momentum and spin. The sign flip of $\mathbf{k}$ can be compensated by applying the time-translation operator $\hat{\mathcal{T}}_{T/2}=\exp\left(i\frac{T}{2}\partial_{t}\right)$, which shifts time by half a period and thereby reverses the sign of the vector potential. Consequently, the combined operator $\hat{g}\equiv\hat{\mathcal{C}}_{2}\otimes\hat{\mathcal{T}}_{T/2}$ serves as a dynamical symmetry of the time-dependent Hamiltonian. In general, a dynamical symmetry $\hat{g}\equiv\hat{\mathcal{P}}\otimes\hat{\mathcal{T}}_{\tau}$ that includes a time translation by a generic interval $\tau$ imposes the following constraint on the $n$-order current
$P{\bf J}_{n}=e^{i\Omega\tau}{\bf J}_{n}^{(*)},$ (2)
where $P$ is the $2\times 2$ matrix associated with the action of $\hat{\mathcal{P}}$ on the two components of the current. Eq. (2) is valid when $\hat{\mathcal{P}}$ is (anti-)unitary, and its derivation is discussed in [28]. Here, we focus on its consequences. For $\hat{\mathcal{C}}_{2}$-symmetric materials illuminated by monochromatic light, whether linearly or circularly polarized, $\tau=T/2$ represents half the period and $P=-\mathbb{1}$ corresponds to minus the 2x2 identity. As a result,
$\hat{H}_{0}\text{ is }\hat{\mathcal{C}}_{2}\text{-invariant}\quad\Longrightarrow\quad{\bf J}_{n}=(-1)^{n+1}{\bf J}_{n}.$ (3)
Eq. (3) leads to the crucial consequence that, for even $n$, ${\bf J}_{n}$ vanishes. Beside, we recall that, as a consequence of the $\hat{\mathcal{C}}_{2}$-invariance, the spin texture $\boldsymbol{\sigma}_{\bf k}\equiv<\hat{\boldsymbol{\sigma}}>_{\bf k}$ transforms as follows:
$\hat{H}_{0}\text{ is }\hat{\mathcal{C}}_{2}\text{-invariant}\quad\Longrightarrow\quad\begin{cases}\sigma_{-{\bf k}}^{xy}=-\sigma_{{\bf k}}^{xy}\\ \sigma_{-{\bf k}}^{z}=\sigma_{{\bf k}}^{z}\end{cases}.$ (4)
Eqs. (3) and (4), whose key messages are illustrated in Fig. (1), constitute the first central result of this paper: in a 2D non-centrosymmetric system where the spin texture is $\hat{\mathcal{C}}_{2}$-invariant the emission of even-order harmonics is forbidden. This important criterion implies that systems with $\hat{\mathcal{C}}_{2}$-broken spin texture can, in principle, emit even-order harmonics. Before discussing this class of materials, we note that the one-to-one correspondence between the HHG selection rules and the spin texture in the presence of SOC holds a more general validity. To provide another example, we consider systems with a generic rotational invariance $\hat{\mathcal{C}}_{N}$ with respect to the angle $\phi=2\pi/N$, where $N=2,3,4,6$, under a circularly-polarized pump field. In this case, the dynamical symmetry is $\hat{g}\equiv\hat{\mathcal{C}}_{N}\otimes\hat{\mathcal{T}}_{T/N}$, and as discussed in detail in [28], it imposes the constraint $Q_{N}{\bf J}_{n}=\exp\left(i2\pi/N\right){\bf J}_{n}$, where $Q_{N}$ is the matrix associated with the rotation of $2\pi/N$. Apart from the phase factor arising from the time-dependence of the driven Hamiltonian, the last equation closely resembles the corresponding constraint on the spin texture, which is $Q_{N}\boldsymbol{\sigma}_{\bf k}=\boldsymbol{\sigma}_{Q_{N}{\bf k}}$. These equations once again highlight the close connection between spin textures and high-order harmonics. When considering vertical mirrors, they impose constraints on the polarization of the harmonics. Here, we focus on a representative result for a pump field that is linearly polarized perpendicular to a vertical mirror, while referring to [28] for a comprehensive analysis of the entire set of cases. Denoting ${\bf J}^{\parallel}$ and ${\bf J}^{\perp}$ as the components of the optical current parallel and perpendicular, respectively, to the mirror direction, we have
$\begin{cases}{\bf J}_{n}^{\parallel}={\bf 0},\quad n\text{ even},\\ {\bf J}_{n}^{\perp}={\bf 0},\quad n\text{ odd}.\end{cases}$ (5)
This can be compared with the corresponding transformation properties of the spin texture, namely, $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\parallel}=-\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\parallel}$ and $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\perp}=\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\perp}$. From this point onward, we focus on even-order harmonics. Having identified the $\hat{\mathcal{C}}_{2}$-invariant spin texture as a sufficient condition for their suppression, we now examine scenarios where spin textures break this symmetry, enabling their emission. In this context, it is essential to distinguish between the two cases in which TR symmetry is either preserved or broken.
Figure 2: Spin textures and high-harmonic response for different physical configurations. Schematic illustration of the trigonal lattice (a) with threefold rotational symmetry and antiferromagnetic pattern (d) with broken twofold rotational symmetry. Both systems possess a vertical mirror symmetry $\mathcal{M}_{x}$ along the $x$-axis, and in both cases, the incident field (green arrow) is polarized perpendicular to this direction. Blue and orange arrows represent the expected polarizations for odd- and even-order harmonics, respectively. Panels (b) and (e) show the spin textures in momentum space corresponding to the physical configuration in (a) and (d), respectively. The color legend indicates the out-of-plane spin component $\langle\sigma_{z}\rangle$, with color blue, gray and red corresponding to $\langle\sigma_{z}\rangle=-1$, 00 and $1$, respectively. Black circles mark spots in the spin textures where $\hat{\mathcal{C}}_{2}$ symmetry, as described by Eq. (4), is broken. HHG spectra for the time-reversal invariant trigonal configuration (c) and for the antiferromagnetic pattern (f), respectively. The Hamiltonian parameters, in units of the hopping energy $t$, are: $\gamma_{R}=0.1$ for both cases; $\lambda=0.1$ in (c); and $\Delta\varepsilon=0.5$, $B=0.2$ in (f). The computations were performed at temperature $T=0$ with chemical potential $\mu=1.0$. In both cases, the incident field has intensity $|{\bf A}_{0}|=1.0$ and frequency $\Omega_{in}=0.5$. Polarizations of odd- (blue dots) and even-order (orange squares, with different shades of orange) harmonics are consistent with the outcomes of Eq. (5). Insets show the intensity of even-order harmonics as functions of the parameters controlling the $\hat{\mathcal{C}}_{2}$-symmetry breaking: the trigonal warping $\lambda$ for (c), and the charge imbalance $\Delta\varepsilon$ for (f).
Systems with broken $\hat{\mathcal{C}}_{2}$ and preserved TR symmetry We consider the spin texture and high-harmonic response of a TR-invariant trigonal lattice with threefold rotational symmetry ($\hat{\mathcal{C}}_{3}$) that explicitly breaks $\hat{\mathcal{C}}_{2}$ symmetry. The Hamiltonian is given by $\hat{H}_{0}=\varepsilon_{\bf k}\hat{\sigma}_{0}+{\bf h}_{\bf k}\cdot\hat{\boldsymbol{\sigma}}$, where $\hat{\sigma}_{0}$ is the identity matrix and $\hat{\boldsymbol{\sigma}}$ represents the Pauli matrices. Here, having defined ${\bf a}_{1/2}=\pm\left(\sqrt{3}/2\right)\hat{{\bf x}}-\left(1/2\right)\hat{{\bf y}}$ and ${\bf a}_{3}=\hat{{\bf y}}$, $\varepsilon_{\bf k}=-2t\left[\cos\left({\bf a}_{1}\cdot{\bf k}\right)+\cos\left({\bf a}_{2}\cdot{\bf k}\right)+\cos\left({\bf a}_{3}\cdot{\bf k}\right)\right]$ is the tight-binding energy dispersion, $h_{\bf k}^{(x)}=-\gamma_{R}\boldsymbol{\eta}_{\bf k}\cdot\hat{{\bf y}}$ and $h_{\bf k}^{(y)}=\gamma_{R}\boldsymbol{\eta}_{\bf k}\cdot\hat{{\bf x}}$, with $\boldsymbol{\eta}_{\bf k}=2\left[{\bf a}_{1}\sin\left({\bf a}_{1}\cdot{\bf k}\right)+{\bf a}_{2}\sin\left({\bf a}_{2}\cdot{\bf k}\right)+{\bf a}_{3}\sin\left({\bf a}_{3}\cdot{\bf k}\right)\right]$, describe Rashba spin-momentum locking, and $h_{\bf k}^{(z)}=2\lambda\left[\sin\left({\bf a}_{1}\cdot{\bf k}\right)+\sin\left({\bf a}_{2}\cdot{\bf k}\right)+\sin\left({\bf a}_{3}\cdot{\bf k}\right)\right]$ is the trigonal warping term. While the Rashba coupling is invariant under any rotation, $h_{\bf k}^{(z)}$ breaks $\hat{\mathcal{C}}_{2}$, reflecting the $\hat{\mathcal{C}}_{3}$ symmetry of the underlying lattice, and preserves a vertical mirror along the $x$ axis, see Fig. (2)(a). It also generates a non-zero Berry curvature, ${\boldsymbol{\Omega}}({\bf k})\propto\lambda$[7]. The spin texture and the HHG spectrum for this system are presented in panels (b) and (c) of Fig.(2), respectively. Here, the HHG spectrum is computed under a monochromatic light pulse linearly-polarized along the $y$ axis, which is orthogonal to the mirror axis. Even-order harmonics exhibit finite amplitude, consistent with the expectation from the spin texture shown in Fig. (2)(b), which breaks $\hat{\mathcal{C}}_{2}$. These harmonics eventually vanish when the $\hat{\mathcal{C}}_{2}$-breaking warping term approaches zero at $\lambda=0$ (see panel (c) and inset therein). Additionally, we observe that odd- and even-order harmonics are polarized perpendicular and parallel to the mirror axis, i.e., respectively, along the $y$ and $x$ axes. This is in agreement with the HHG selection rules in Eq. (5) for a pump field oriented perpendicular to a vertical mirror. It is relevant now to consider the fate of even-order harmonics in a system with broken $\hat{\mathcal{C}}_{2}$ symmetry but zero Berry curvature. An example of such a system is one characterized by Rashba coupling and $h_{\bf k}^{(z)}=\zeta k_{x}$, where ${\boldsymbol{\Omega}}({\bf k})={\bf 0}$ throughout momentum space. Explicit calculations demonstrate that, in this case, even-order harmonics vanish. This serves as a particular example of a more general result, as shown in [28]: in TR-invariant systems where the Berry curvature is zero throughout the momentum space, the emission of even-order harmonics is forbidden. Therefore, we conclude that, in 2D non-centrosymmetric materials with TR symmetry, a spin texture that breaks $\hat{\mathcal{C}}_{2}$ symmetry, combined with a finite Berry curvature, ensures the non-vanishing of even-order harmonics.
Systems with broken $\hat{\mathcal{C}}_{2}$ and T-symmetry — To show that even-order harmonics can effectively probe $\hat{\mathcal{C}}_{2}$-breaking magnetic patterns in systems with broken inversion and TR symmetry, we examine the case of a square crystal that is intrinsically invariant under $\hat{\mathcal{C}}_{2}$ symmetry, but exhibits an antiferromagnetic spin arrangement and a charge imbalance (Fig. 2(d)) within the sub-lattice that together break the twofold rotational symmetry. The Hamiltonian reads $\hat{H}_{0}=\varepsilon_{\bf k}^{\mu}\hat{\tau}_{\mu}\otimes\hat{\sigma}_{0}+h_{\bf k}^{\mu i}\hat{\tau}_{\mu}\otimes\hat{\sigma}_{i}$, where $\hat{\tau}_{\mu}$ are Pauli matrices representing sub-lattice degrees of freedom. To ensure $\hat{\mathcal{C}}_{2}$ symmetry breaking, we assume a finite charge imbalance $\Delta\varepsilon$ within the sub-lattice, $\varepsilon_{\bf k}^{z}=\Delta\varepsilon$, and an antiferromagnetic pattern with a local magnetic field $B$ oriented along the $\hat{y}$ direction, $h_{\bf k}^{zy}=B$, as illustrated in Fig. 2(d). We also include Rashba coupling, $h_{\bf k}^{xx}=-\gamma_{R}\sin k_{y}$ and $h_{\bf k}^{yx}=\gamma_{R}\sin k_{x}$. The resulting spin texture breaks rotational invariance, as clear from Fig. 2(e), and even-order harmonics become non-zero, as demonstrated in Fig. 2(f). These harmonics eventually vanish in the limit of zero charge imbalance, as indicated in the inset of Fig. 2(f), since the antiferromagnetic pattern alone is $\hat{\mathcal{C}}_{2}$-invariant. As a concluding remark, we address the role of Berry curvature in enabling even-order harmonics when TR symmetry is broken. In this case, the spin texture may exhibit invariance under $\hat{\mathcal{C}}_{2}$, thereby ensuring the vanishing of even-order harmonics, while the Berry curvature remains finite. An example of this scenario is a 2D ferromagnet with Rashba coupling and spin alignment along the out-of-plane direction. This case illustrates that, in the absence of TR symmetry, the rotational symmetry of the spin texture is the key factor in the emission of even-order harmonics.
Conclusions — We have investigated the interplay between spin textures, Berry curvature, and HHG in two-dimensional non-centrosymmetric systems. Our analysis reveals that the presence or absence of even-order harmonics is closely tied to the symmetries of the spin texture. Specifically, we established that spin textures with broken twofold rotational symmetry are essential for the emission of even-order harmonics. Furthermore, we demonstrated the crucial role of Berry curvature in enabling even-order harmonics in systems with time-reversal symmetry. We emphasize that rotational symmetry can be broken across various degrees of freedom, all linked to the high-harmonic response through spin-orbit interaction. This means that even-order harmonics can identify not only crystalline and spin patterns lacking rotational symmetry but also emergent electronic orderings with spatial reconstructions, such as in kagome systems, where trimerization breaks six-fold rotational symmetry[32]. Moreover, nontrivial high-harmonic response can also occur in systems exhibiting specific patterns of spin [33, 34] and orbitally polarized currents. Our results can also be applied to investigate spin ordering in low-dimensional altermagnets [35], marked, for specific terminations, by surface spin textures with odd-parity multipole patterns that break twofold rotational symmetry. All these rotational symmetry breakings, and the associated phase transitions, can be observed through the shift from zero to finite even-order harmonic amplitudes, or conversely, when twofold symmetry is restored from a lower symmetry state.
Acknowledgements
F.F. and F.G. acknowledge support from the Italian Ministry of University and Research (MUR), under Grant PRIN No. 2020JZ5N9M (CONQUEST) and Grant PRIN No. 2022TWZ9NR (STIMO). M.C. acknowledges partial support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE) and, together with F.F., from PNRR MUR project PE0000023-NQSTI. C.O. acknowledges support by the Italian Ministry of Foreign Affairs and International Cooperation, Grant No. PGR12351 (UTLRAQMAT) and the PNRR MUR project PE0000023-NQSTI (TOPQIN).